### v1

# https://zielon.github.io/gem/

Abstract
We present personalized Gaussian Eigen Models (GEMs) for human heads, a novel method that compresses dynamic 3D Gaussians into low-dimensional linear spaces. Our approach is inspired by the seminal work of Blanz and Vetter, where a mesh-based 3D morphable model (3DMM) is constructed from registered meshes. Based on dynamic 3D Gaussians, we create a lower-dimensional representation of primitives that applies to most 3DGS head avatars. Specifically, we propose a universal method to distill the appearance of a mesh-controlled UNet Gaussian avatar using an ensemble of linear eigenbasis. We replace heavy CNN-based architectures with a single linear layer improving speed and enabling a range of real-time downstream applications. To create a particular facial expression, one simply needs to perform a dot product between the eigen coefficients and the distilled basis. This efficient method removes the requirement for an input mesh during testing, enhancing simplicity and speed in expression generation. This process is highly efficient and supports real-time rendering on everyday devices, leveraging the effectiveness of standard Gaussian Splatting. In addition, we demonstrate how the GEM can be controlled using a ResNet-based regression architecture. We show and compare self-reenactment and cross-person reenactment to state-of-the-art 3D avatar methods, demonstrating higher quality and better control. A real-time demo showcases the applicability of the GEM representation.
Keywords:
Eigen Head Avatars, Gaussian Splatting
Figure 1: Given a multi-view video of a subject and meshes in full correspondence, we construct a high-quality Gaussian Eigen Model (GEM). Our method, employing a UNet (left), regresses a sequence of 3D Gaussians that undergo distillation to obtain compressed linear eigenbasis which we call Eigen Gaussians (right). To generate new photo-realistic 3D faces only a linear combination of the Gaussian basis has to be computed. We show that the required coefficients for the linear combination can be regressed from single images, allowing for real-time facial animation. The simplicity of a GEM leads to very efficient storage and rendering times.
1 Introduction
Half a century ago, Frederick Parke described a representation and animation technique to generate ,,animated sequences of a human face changing expressions” [48]. Using polygonal meshes, single facial expression states were described that could be combined with linear interpolation to generate new expression states (the ,,simplest way consistent with natural motion” [48]). Based on this principle, Blanz, and Vetter [3] introduced the so-called 3D morphable model (3DMM) - a statistical model of the 3D shape and appearance of human faces. Principle Component Analysis (PCA) is performed on a set of around 200 subjects that have been laser scanned and registered to a consistent template to find the displacement vectors (principal components) of how faces change the most, in terms of geometry and albedo. With this PCA basis, new faces can be generated by specifying the coefficients for the principle components taking a dot product of the coefficients with the basis to obtain offests, and adding them to the mean. State-of-the-art reports on face reconstruction and tracking [86], as well as on morphable models [8], state that this representation is widely used for performance capturing (regression-based and optimization-based) and builds the backbone of recent controllable photo-realistic 3D avatars that are equipped with neural rendering [63, 64, 81, 17, 11].
Inspired by the simplicity of such mesh-based linear morphable models and addressing the lack of appearance realism of current 3DMMs, we propose a personalized linear appearance model based on 3D Gaussians as geometry primitive following 3D Gaussian Splatting (3DGS) [28]. In contrast to the concurrent work on Dynamic 3D Gaussian Avatars [38, 53, 83, 47, 79, 74, 59], our goal is a compact representation that does not need vast amounts of compute resources to generate novel expressions of the human. Unfortunately, most of the methods show that to produce high-quality results, one needs to employ heavy CNN-based architectures which are not well suited for commodity devices and tend to slow down the rendering pipeline. Moreover, those models comprise dozens of millions of parameters creating heavy checkpoints that can easily exceed 500 MB. This ultimately creates a major issue for distributing and managing personalized models. We tackle this problem by distilling a novel CNN-based architecture, leading to a personalized Gaussian Eigen Model, GEM in short. Our approach builds on Gaussian maps predicted from a modified UNet architecture [69] which is used for the UV space normalization required to properly build linear eigenbasis. Based on the per-subject trained model, we bootstrap the personalized GEM by computing an ensemble of linear basis on the predicted Gaussian maps of the training frames. Next, the basis is refined on the training corpus using photometric losses while preserving the orthonormality.
This lightweight appearance basis is controlled with a relatively low number of parameters ranging from twenty up to fifty coefficients which can be specified w.r.t. the available compute resources and can for example be regressed by a ResNet-based model [9]. We demonstrate this for the scenarios of self-reenactment as well as cross-person animation. We showcase a real-time application of our method in the supplemental video.
In summary, our contributions are:
  1. 1.
    Gaussian Eigen Model (GEM) a universal distillation technique of 3D Gaussian head avatar model that is built upon an ensemble of eigenbasis.
  2. 2.
    A state-of-the-art CNN-based avatar appearance model that can be constructed using multi-view captures of a subject and serves as an example that can be used in combination with GEM.
  3. 3.
    Real-time (cross-person) animation of GEMs from single input images using a generalizable regressor.
2 Related Work
The majority of current face representations and tracking techniques are based on parametric 3D morphable models (3DMM) [3, 37]. For a detailed overview of 3DMM-based reconstruction and appearance models, we refer to the state-of-the-art reports on face tracking and reconstruction [86], the report on morphable models [8], and the two neural rendering state-of-the-art reports [63, 64] that demonstrate how neural rendering can be leveraged for photo-realisitic facial or full body avatars. In the following, we review the recent methods that are focused on photo-realistic 3D avatars which build appearance models using neural radiance fields (NeRF) [43] or volumetric primitives like 3D Gaussians [28].
2.1 NeRF-based avatars
One of the first methods that combined 3DMM and NeRF is NeRFace [11]. In this work, a neural radiance field is directly conditioned by expression codes using the Basel Face Model (BFM) [3] optimized by the Face2Face tracker [67]. This idea gave rise to many methods [84, 81, 17, 82, 78, 76, 52, 12] following a similar approach, but attaching the radiance fields more explicitly to the surface of the 3DMM, e.g., by using the 3DMM-defined deformation field. To enable photorealistic results some methods employ StyleGAN2 [27] with a NeRF-based renderer [26, 1, 5]. Generative methods like EG3D [5] and PanoHead [1] employ a GAN-based training to predict triplane features that span a neural radiance field. GANAvatar [26] applies this scheme to reconstruct a personalized avatar.
Close to our method is StyleAvatar [69]. Based on 3DMM tracking the method learns a personalized avatar that benefits from a StyleUNet which incorporates StyleGAN [27] to decode the final image. Despite real-time capabilities, the method suffers from artifacts produced by the image-to-image translation network that we explicitly avoid by using Gaussian maps which can compensate for tracking misalignments by predicting corrective fields for the 3D Gaussians.
2.2 3D Avatars from Volumetric Primitives
Neural Volumes (NV) [39] started an impressive line of research in the context of digital avatars. Using multiview images with a Variational Auto-Encoder [30] and volumetric integration it encodes arbitrary dynamic scenes into a volume which can be deformed by traversing the latent code z. To better control the 3D space, Lombardi et al. [40] introduced Mixture of Volumetric Primitives (MVP) a hybrid representation based on primitives attached to the tracked mesh which ultimately replaced the encoder from NV. Each primitive is a volume represented as a small voxel with $32^{3}$ cells that on the vertices store RGB and opacity values. The final color is obtained by integrating values along a pixel ray. This hybrid representation became very popular spawning many follow-up projects [4, 55, 36, 62, 71]. The recently presented 3D Gaussians Splatting (3DGS) [28] introduced an alternative representation to MVP primitives. 3DGS represents a volume as a set of anisotropic 3D Gaussians, which are equivalently described as ellipsoids, in contrast to isotropic spheres used in Pulsar [35]. Similarly to DSS [70] it defines a backward pass of the rasterization technique Surface Splatting introduced by Zwicker et al. [87]. However, with the difference of using $\alpha$-blending [43] to integrate 3D Gaussians along a given ray. Numerous recent methods [83, 38, 59, 74, 80, 53, 10, 73, 24, 47, 32, 15] capitalize on the speed and quality of 3DGS. Qian et al. attached 3D Gaussians to the FLAME [37] mesh surface and applied a deformation gradient similar to Zielonka et al. [84] to orient the Gaussians according to the local Frenet frames of the surface. This method, however, does not utilize any information about expressions and, thus, struggles with pose-dependent changes (e.g., wrinkles, self-shadows) and, despite high-quality results, retrieves only a global static appearance model. Li et al. [38] are using the StyleUNet architecture [69] to regress front and back Gaussian maps. Employing the powerful CNN network on position maps, they achieve impressive results for human bodies. However, their architecture uses three separate StyleUNets each with two decoders for front and back texture which creates significant computational overhead making the whole pipeline slow. In contrast, we propose a more efficient CNN-based model that operates in a UV-space which we distill a lightweight linear 3D Gaussian Eigen Model (GEM).
3 Method
Current dynamic 3D Gaussian Avatars show unprecedented quality, however, they require sophisticated and often compute-heavy CNN-based architectures [38, 47, 74] leveraging large receptive fields to capture high-frequency details like pose-dependent wrinkles or self-shadows. We present an efficient CNN-based high-quality model and propose a technique to compress and distill such an avatar into an ensemble of eigenbasis that preserves its neural counterpart’s quality and expressiveness. The following section will explain our novel CNN-based appearance model and how to distill it into an alternative representation called the Gaussian Eigen Model (GEM).
Figure 2: CNN-based Gaussian Avatar Pipeline. Our CNN model produces delta Gaussian maps $\boldsymbol{D}_{i}$ [38, 47] and static color $\boldsymbol{C}$ from a multi-view video. Similarly to Animatable Gaussians [38], we constrain the network to operate in a reduced linear space, i.e., the mesh $\boldsymbol{M}_{i}$ is projected on a PCA basis $\boldsymbol{M}_{PCA_{i}}=\Gamma_{\text{PCA}}(\boldsymbol{M}_{i})$ which is then input to the network $\Pi$ after converting the mesh to a normal map $\phi(\boldsymbol{M}_{PCA_{i}})$.
Note that compared to other neural avatars [84, 11, 17, 39, 74, 69], GEM does not require a 3DMM [37, 50], i.e., training GEM only requires meshes in correspondence, while testing uses the GEM model parameters, making it effectively mesh and neural network free.
3.1 3D Gaussian Splatting Preliminaries
3D Gaussian Splatting (3DGS) [28] is an alternative approach to Neural Radiance Field (NeRF) [43] for static multi-view scene reconstruction and rendering under novel view. Kerbl et al. [28] parameterize the space as scaled 3D Gaussians [70, 34] with a 3D covariance matrix $\mathbf{\Sigma}$ and mean $\mathbf{\mu}$:
$G(\mathbf{x})=e^{-\frac{1}{2}(\mathbf{x-\mu})^{T}\mathbf{\Sigma}^{-1}(\mathbf{x-\mu})}.$ (1)
To render this representation, Zwicker et al. [87] employ the projection of 3D Gaussians onto the image plane using the formula $\mathbf{\Sigma}^{\prime}=\mathbf{A}\mathbf{W}\mathbf{\Sigma}\mathbf{W}^{T}\mathbf{A}^{T}$, where $\mathbf{\Sigma}^{\prime}$ represents the covariance matrix in 2D space. Here, $\mathbf{W}$ denotes the view transformation, and $\mathbf{A}$ represents the projective transformation. To avoid direct optimization of the covariance matrix $\mathbf{\Sigma}$ which must be positive semidefinite, Kerbl et al. [28] use scale $\mathbf{S}$ and rotation $\mathbf{R}$ which equivalently describes 3D Gaussian as a 3D ellipsoid $\mathbf{\Sigma}=\mathbf{R}\mathbf{S}\mathbf{S}^{T}\mathbf{R}^{T}$. Finally, 3DGS follows Ramamoorthi et al. [54] to approximate the diffuse part of the BRDF [16] as spherical harmonics (SH) to model global illumination and view-dependent color. Four bands of SH are used which results in a 48 elements vector.
3.2 Appearance Maps Generator
As an example of a powerful Gaussian generator, we present a CNN-based appearance model inspired by Pix2Pix [22], which translates an input image to a specific output domain. In our case, Gaussian maps are generated from coarse normal maps of a tracked mesh. The idea of organizing the 3D Gaussians in 2D maps is used in several methods  [38, 47, 74]; each texel in such a map represents a 3D Gaussian with its parameters: offset vector for rotation, scale, position, and opacity. Similarly to StyleAvatar [69], we use a StyleGAN-based [27] encoder and decoder network for this image translation. However, in contrast to StyleAvatar, we propose a more lightweight image translation pipeline, where we reduce the number of encoders from three to two and the number of decoders from six to two, and decrease the size of the StyleGAN [27] decoder. To efficiently use the 2D map space, we do not use projective textures from meshes [69], but a UV parametrization which reduces half of the decoders in comparison to StyleAvatar. Moreover, we do not use triplets of the encoder-decoder, as we combine all properties of the Gaussians into one map $\boldsymbol{G}_{i}$ following ASH [47]. Therefore, we define our network as follows:
$\begin{split}\boldsymbol{D}_{i}&:\Pi(\phi(\Gamma_{\text{PCA}}(\boldsymbol{M}_{i}))),\\ \boldsymbol{C}&:\Psi(\phi(\boldsymbol{M})),\\ \end{split}$ (2)
where $\boldsymbol{D}_{i}\in\mathbb{R}^{11\times H\times W}$ is a map containing the delta for positions $\Delta_{pos}\in\mathbb{R}^{3}$, rotation $\Delta_{rot}\in\mathbb{R}^{4}$, scale $\Delta_{scale}\in\mathbb{R}^{3}$, and opacity $\Delta_{op}\in\mathbb{R}$. The colors $\boldsymbol{C}\in\mathbb{R}^{3\times H\times W}$ of the Gaussians are predicted using the normal map $\phi(\boldsymbol{M})$ of the canonical mesh $\boldsymbol{M}$ ($\phi$ is the normal map extractor from a mesh). Similar to Animatable Gaussians [38], we use a PCA layer $\Gamma_{\text{PCA}}$ which serves as a low-pass regularization filter for the input. $\Gamma_{\text{PCA}}$ is built by using PCA on the meshes $\boldsymbol{M}_{i}$ for the training frames and 16 principle components are used as the basis. For the training, we use the projection of each incoming mesh $\boldsymbol{M}_{i}$ on the PCA manifold $\boldsymbol{M}_{PCA_{i}}=\Gamma_{\text{PCA}}(\boldsymbol{M}_{i})$.
The final Gaussian map $\boldsymbol{G}_{i}\in\mathbb{R}^{11\times H\times W}$ is obtained by applying the deltas to the canonical Gaussians $\boldsymbol{G}$ [38, 83]. The deformed position of 3D means is computed as $\boldsymbol{T}_{i}(\boldsymbol{M}+\boldsymbol{D}_{i_{0:3}})$ (Section 3.3). Note that our color map C is static and does not model view-dependent effects; this means that we force the network to recover globally consistent colors for each Gaussian similar to a texture in the classic 3DMM. Therefore, $\boldsymbol{G}_{i}$ must model the wrinkles and self-shadows.
Finally, we use Gaussian splatting [28] to render the regressed Gaussian maps. We define the predicted color of pixel $(u,v)$ as:
$\mathbf{\bar{C}}_{u,v}=\sum_{i\in\mathcal{N}}\mathbf{c_{i}}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}),$ (3)
where $\mathbf{c}_{i}$ is the Gaussian color predicted by $\Psi$, $\mathcal{N}$ is the number of texels and $\alpha_{i}$ is predicted opacity per Gaussian.
3.3 Deformation Model
The output of the CNN regressors $\Pi$ and $\Psi$ are in a canonical space. For the transformation from the canonical space to the deformed space, we employ deformation gradients, following Sumner et al. [60]. This approach allows for greater flexibility regarding input meshes, provided they maintain full correspondence. Given a mesh $\boldsymbol{M}_{PCA_{i}}$ for the frame $i$, we define the deformation gradients as $\mathbf{J}_{j}=\hat{\mathbf{E}}_{j}\mathbf{E}_{j}^{-1}$, where $\hat{\mathbf{E}}_{j}\in\mathbb{R}^{3\times 3}$ and $\mathbf{E}_{j}\in\mathbb{R}^{3\times 3}$ contain the Frenet frame (tangent, bi-tangent, normal) of the triangle $j$ defined in deformed and canonical spaces, respectively. Using these deformation gradients and the known correspondences between the Gaussian map and the meshes, we transform the Gaussians from the canonical space to the deformed space.
3.4 Implementation Details
The training objective of the CNN-based appearance model is defined as $\mathcal{L}=\mathcal{L}_{Color}+\mathcal{L}_{Reg}$. $\mathcal{L}_{Color}$ is a weighted sum of three different photo-metric losses between the rendered image $\mathbf{\bar{C}}$ and the ground truth $\mathbf{C}$:
$\begin{split}\mathcal{L}_{Color}&=(1-\omega)\mathcal{L}_{1}+\omega\mathcal{L}_{\text{D-SSIM}}+\zeta\mathcal{L}_{\text{VGG}},\\ \mathcal{L}_{Reg}&=\lambda\sum_{j=1}^{N}\left\lVert\Delta_{pos_{j}}\right\rVert^{2}+\gamma\sum_{j=1}^{N}\left\lVert s_{scale_{j}}\right\rVert^{2},\\ \end{split}$ (4)
where $\omega=0.2$, $\zeta=0.0075$ (after 150k iterations steps and zero otherwise), $\mathcal{L}_{\text{D-SSIM}}$ is a structural dissimilarity loss, and $\mathcal{L}_{\text{VGG}}$ is the perceptual VGG loss. $\mathcal{L}_{Reg}$ regularizes position offsets $\Delta_{pos_{j}}$ and scales $s_{scale_{j}}$ to stay small w.r.t. the input mesh. We train our model for $10^{6}$ steps using Adam [29] with a learning rate $5\mathrm{e}{-4}$ and a batch size of one which takes around 24h on a Nvidia A100.
Our method and all the baselines were trained using the same multiview input data sourced from the dataset provided by Qian [53], which includes multiview images from the NeRSamble dataset [33] as well as tracked meshes. These datasets are licensed and managed by MPI and TuD.
4 Gaussian Eigen Model (GEM)
Once a powerful appearance generator is available, we can build our universal eigenbasis model, GEM. For each training frame, we save the predicted Gaussian maps $\boldsymbol{G}_{i}$, where the positions are normalized by un-posing the head joint. The Gaussian maps could be successfully replaced with Gaussian point clouds, as long as their quantity remains unchanged. We stack the Gaussian maps in a matrix $\mathbf{A}\in\mathbb{R}^{N\times 11\times H\times W}$ where $N$ is the number of frames. Since the color $\boldsymbol{C}$ was optimized globally, it acts like a classic texture and therefore does not need distillation.
Figure 3: Samples of a GEM. We display samples for the first three components of the geometry eigenbasis of a GEM in the range of $[-3\sigma,3\sigma]$, showing diverse expressions. Note that GEM requires no parametric 3D face model like FLAME[37].
We build our statistical model for each modality of the Gaussians from $\mathbf{A}$ separately, specifically, we have individual bases for rotation, position, opacity, and scales. To compute the ensemble of eigenbasis we use Scikit-learn [51] and PCA [25] and define our model as an orthogonal coordinate system formed by the eigenvectors of the covariance matrices of each modality, formally we solve:
$(\mathbf{A}^{T}\mathbf{A})\mathbf{V}=\mathbf{V}\mathbf{\Lambda},$ (5)
where $\mathbf{A}^{T}\mathbf{A}\in\mathbb{R}^{M\times M}$ is the covariance matrix, $\mathbf{V}\in\mathbb{R}^{M\times M}$ is the matrix of eigenvectors and $\mathbf{\Lambda}\in\mathbb{R}^{M\times M}$ is the diagonal matrix of eigenvalues ($M=N\times 11\times H\times W$). Thus, per modality, we have:
$\mathbf{S}_{mod}=\mathbf{\bar{S}}_{mod}+\sum^{m}_{i=1}\lambda_{i}\mathbf{v}_{i},$ (6)
where $\lambda_{i}$ are coefficient ($\mathbf{\Lambda}_{:,i}$), $\mathbf{\bar{S}}_{mod}$ is a given modality mean, $\mathbf{v}_{i}$ are the spanning vectors of the corresponding orthonormal basis $\mathbf{V}_{:,i}$, and the number of principal components is set to $m=50$. 3 depicts samples of a GEM on the 3D position modality. For the remainder of the document, we omit the mod index for simplicity.
After estimating the initial basis $\mathbf{V}$ which might contain artifacts produced by the CNN network, we refine it using a photometric reconstruction loss on the training images for around 30k iterations employing the same objectives from Equation 4 to obtain $\mathbf{\hat{V}}$. To ensure that the basis stays orthonormal throughout this refinement, every 1k steps, we orthogonalize $\mathbf{\hat{V}}$ using QR decomposition such that $\mathbf{\hat{V}}^{T}\mathbf{\hat{V}}=\mathbf{\hat{V}}\mathbf{\hat{V}}^{T}=\mathbf{I}$. Example training PSNR errors w.r.t refinement steps 0k, 5k, and 30k are 34.75dB, 36.68dB, and 36.85dB, respectively, showing further improvement in the quality.
4.1 Image-based animation of GEM
Expressions for a GEM are fully defined by their coefficients. This is a similar idea to codec avatars introduced by Ma et al. [42], however, our approach does not need additional small regression MLPs. There are several ways to obtain the coefficients of a GEM, for example, one could employ analysis-by-synthesis-based optimization or regression. Analysis-by-synthesis is the backbone of current avatar methods, as they use photometric or depth-based face trackers to sequentially optimize the coefficients of the underlying 3DMMs like FLAME [67, 84, 17] which is typically slow. As a fast, but more imprecise alternative, regressors like DECA [9] or EMOCA [7] can be used which are built on a ResNet backbone and regress FLAME parameters directly from an image.
Figure 4: Image-based animation. One of the applications of our GEM is real-time (cross)-reenactment. For that, we utilize generalized features from EMOCA [7] and build a pipeline to regress the PCA coefficients of our model from an input image/video.
As we aim for fast and expressive motion capturing, we build a personalized GEM regressor, that builds upon EMOCA, however, with several major modifications, see 4. First, we do not use FLAME parameters, instead, we use features extracted from a pre-trained EMOCA network denoted as $\Theta(\mathbf{I}_{i})$ where $\mathbf{I}_{i}$ is the current image. EMOCA’s architecture comprises two ResNet networks; one to extract expression features $\mathbf{f}_{expr}\in\mathbb{R}^{2048}$ and the second for shape $\mathbf{f}_{shape}\in\mathbb{R}^{2048}$, both are followed by final MLPs to regress corresponding FLAME parameters. In contrast, we remove the last hidden layer of the final MLP obtaining features $\in\mathbb{R}^{1024}$ and combine the regressed vectors into one $\mathbf{f}_{i}\in\mathbb{R}^{2\times 1024}$ vector. For training, we select five frontal cameras from NeRSemble [33]. Next, we build a PCA layer using the training frames similar to the one from our appearance model, however, with the difference that we use relative features $\mathbf{r}_{i}=\mathbf{f}_{i}-\mathbf{f}_{neutral}$ instead of meshes. The neutral reference frame $\mathbf{f}_{neutral}=\Theta(\mathbf{I}_{neutral})$ is selected manually from the video similar to Face2Face [67]. During training, for each frame, we project $\mathbf{r}_{i}$ onto the PCA manifold to restrict and regularize training using the first 50 components. The projected coefficients are passed through a small MLP that produces a vector of GEM coefficients per modality, denoted as $\boldsymbol{p}_{i}\in\mathbb{R}^{M}$:
$\begin{split}\mathbf{r}_{i}&=\Theta(\mathbf{I}_{i})-\mathbf{f}_{neutral},\\ \boldsymbol{p}_{i}&=3\sigma\cdot\tanh(\textrm{MLP}(\Gamma_{\text{PCA}}(\mathbf{r}_{i}))),\end{split}$ (7)
where $\Gamma_{\text{PCA}}$ is the relative PCA model and $M$ is the selected amount of linear basis. The MLP has three hidden layers with 256 neurons each and ReLU activations. We use a scaled $\tanh$ activation function for the output, to restrict the prediction to be in $[-3\sigma,3\sigma]$. The final 3D model is obtained by $\mathbf{\bar{S}}+\boldsymbol{p}_{i}^{T}\mathbf{\hat{S}}$ which is splatted using 3DGS. We follow the objective function from Equation 4 for supervision on the color reproduction with an additional L1 regularization w.r.t. the corresponding coefficients. We use AdamW with a learning rate $1\mathrm{e}{-3}$ and train it for 200k iterations with batch size one on a single Nvidia Quadro A6000 which takes around 4h.
5 Results
Refer to captionGround TruthOurs GEMOurs NetGA [53]AG [38]INSTA [84]
Figure 5: Novel view synthesis. Both, the network and GEM, show better performance on novel views, especially, in the region of the mouth interior and wrinkles. In this experiment, we are following the evaluation of Gaussian Avatars [53] and demonstrate novel viewpoint generation. GEM is obtained throughout analysis-by-synthesis fitting [3, 67]. Note that the expressions are seen during training.
We evaluate our appearance model against state-of-the-art methods on the NeRSemble [33] dataset. We use the tracked meshes provided by Qian et al. [53] and the synchronized images from 16 cameras with a resolution of $802\times 550$. Our main baselines are Gaussian Avatars (GA[53] which is neural network-free (Gaussians are attached to the FLAME model), our implementation of Animatable Gaussians (AG)  [38] which is based on CNN-predicting Gaussian maps, and INSTA [84] which uses dynamic NeRF [43]. For the comparison, we present both of our appearance models, the StyleUNet-based architecture (Ours Net) and the distilled linear eigen model (Ours GEM) which we evaluate using analysis-by-synthesis fitting to the target images following [85, 67, 3]. Additionally, we show cross-reenactment results based on relative FLAME parameters transfer [84] as well as our coefficients regressor versus other baselines which are using FLAME meshes regressed by DECA [9] instead of the ground truth ones from [53]. All of the methods are evaluated using several image space metrics on novel expressions and novel views, following the test and novel-view split from Qian et al. [53]. For all of our experiments with eigenbasis, we use 50 components distilled from $256^{2}$ textures which give around 60k active Gaussians. Animatable Gaussians [38] use a similar amount of primitives for front and back textures and Gaussian Avatars [53] typically use around 100k Gaussians.
Table 1: Evaluation on novel expressions and views shows improved results of our GEM optimized using analysis-by-synthesis compared to others. See 6 for the corresponding qualitative results.
Method PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$ L1 $\downarrow$
Ours GEM 32.6781 0.0675 0.9633 0.0069
Ours Net 29.2454 0.0777 0.9448 0.0096
AG [38] 29.0114 0.0812 0.9429 0.0099
GA [53] 28.3137 0.0815 0.9433 0.0102
INSTA [84] 27.9181 0.1153 0.9340 0.0128
Table 2: Novel viewpoint evaluation on a withhold camera, corresponding to Figure 5. Note that the expressions have been seen during training, and only the view is new, the same as in GA [53].
Method PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$ L1 $\downarrow$
Ours GEM 33.5528 0.0678 0.9662 0.0061
Ours Net 32.4622 0.0713 0.9617 0.0067
AG [38] 32.4166 0.0712 0.9614 0.0066
GA [53] 31.3197 0.0786 0.9567 0.0075
INSTA [84] 27.7786 0.1232 0.9294 0.0163
5.1 Image Quality Evaluation
To evaluate our method, we measure the color error in the image space using the following metrics: PSNR (dB), LPIPS [77], L1 loss, and structural similarity (SSIM). We follow the evaluation scheme from Gaussian Avatars [53], using their train and validation split. The evaluation of GEM was generated by sequentially fitting the coefficients to each image using the same objectives from Equation 4. Note that the baselines use the FLAME model with offsets for the tracking, while GEM can directly be used for modeling.
Table 2 presents results on novel expressions evaluated on all 16 cameras. Both the quantitative and qualitative results depicted in Figure 6 show that our PCA model produces fewer artifacts, especially for regions like teeth or facial wrinkles. Table 2 contains an evaluation where we measure errors on novel viewpoints. The results show that our CNN-based appearance model outperforms other neural methods, and, our linear eigenbasis GEM achieves the highest quality due to the ’direct’ analysis-by-synthesis which can fully utilize the expressiveness and grade of our photo-realistic appearance model and is not restricted by 3DMMs like FLAME. Moreover, Figure 5 shows qualitative results of our method on novel views. As can be seen, we better capture high-frequency details, pose-dependent wrinkles, and self-shadows - something which is not possible for methods like Gaussian Avatars [53] or INSTA [84] since they either do not use pose-dependent neural networks or limit the conditioning to a small region only.
Refer to captionGround TruthOurs GEMOurs NetGA [53]AG [38]INSTA [84]
Figure 6: Novel view and expression synthesis. Our Gaussian Eigen Model shows better results in regions like teeth, wrinkles, and self-shadows compared to other methods that struggle with artifacts.
5.2 Cross-person Reenactment Evaluation
Facial cross-person reenactment transfers expressions from the source actor to the target actor. Most of the methods require tracked meshes obtained by fitting the FLAME model for each frame of the source actor sequence. From this tracking, the expression parameters are extracted and the relative change is applied to the neutral target face [67], see Figure 7. While this is an effective approach, it requires having the tracked mesh in a consistent mesh topology, which is a computationally intensive problem, especially in multi-view capture setups.
Table 3: Cross-re-enactment evaluation employing EmoNet features and FID score. *FPS omits the EMOCA forward pass time, which adds approximately 21 FPS to the inference.
Method $\textbf{E}_{feat}$cos $\uparrow$ $\textbf{E}_{feat}$ $\mathcal{L}_{1}$downarrow FID $\downarrow$ *FPS $\uparrow$
AG 0.9396 5.3399 0.4093 16.51
GA 0.8917 6.6141 0.5593 142.71
INSTA 0.9087 6.3153 0.5299 20.62
Ours Net 0.9440 5.1044 0.3685 35.77
Ours GEM 0.9381 5.3197 0.4286 201.70
An alternative way is to use a (monocular) regressor like EMOCA [7] which, in real-time, can predict FLAME parameters for a single frame. We showcase this in comparison to our GEM which is driven by our image-based regressor, see Figure 8. As can be seen in both experiments, our network-based and GEM can produce sharp results. The baseline methods have problems extrapolating to the new expressions, showing severe artifacts in the appearance. Our personalized approach effectively regularizes the regressed coefficients such that the predicted avatar stays in the training distribution effectively avoiding artifacts like in INSTA or Gaussian Avatars.
Drawing inspiration from EMOCA [7], we further assess cross-re-enactment quantitatively by leveraging emotion recognition feature vectors from both the source image and the resulting cross-re-enactment, utilizing EmoNet [68]. For each pair of input and output images, we predict EmoNet features and measure cosine distance and $\mathcal{L}_{1}$ error between them. We report the numbers in the Table 3. Additionally, we also report FID scores [19] and rendering speed. Our method achieves on-par quality with a CNN-based solution while maintaining the highest frame rates and significantly outperforming GA in terms of quality.
Refer to captionGround TruthOurs NetGA. [53]AG. [38]INSTA [84]
Figure 7: Facial cross-person reenactment. The person’s expressions on the left are transferred to the respective avatars on the right. In this experiment, we are using relative expressions based on ground truth meshes from the dataset (FLAME-based meshes reconstructed from multi-view data). Note that this experiment does not apply to our GEM, since it is mesh-free.
Refer to captionGround TruthOurs GEMOurs NetGA [53]AG [38]INSTA [84]
Figure 8: Facial cross-person reenactment using an image-based regressor. The reenactment of the baselines is performed using relative transfer between FLAME meshes regressed by EMOCA compared to our GEM regressor network (Ours GEM).
Table 4: Ablation of GEM. Even with 10 principle components a high PSNR of $31.81dB$ is achieved, while taking only 7MB of memory. In contrast to fixed-sized neural networks, the GEM can be adjusted on the fly depending on the hardware. Moreover, since evaluation requires a single dot product for forward pass the rendering speed is around four times higher than our network. The speed evaluation was done using a single Nvidia A100 GPU.
#Comp $128^{2}$ $256^{2}$ $512^{2}$
PSNR $\uparrow$ Size MB FPS $\uparrow$ PSNR $\uparrow$ Size MB FPS $\uparrow$ PSNR $\uparrow$ Size MB FPS $\uparrow$
10 31.81 7 237.96 31.88 28 210.03 32.23 113 130.46
30 34.20 20 241.31 34.17 83 208.19 34.84 333 112.73
50 34.67 34 238.7 34.61 138 201.70 35.45 553 117.45
Ours Net 33.97 82 47.70 34.99 109 35.77 35.02 178 26.31
AG [38] 33.77 487 18.93 34.40 529 16.51 35.15 636 13.08
5.3 GEM Compression Ablation
Refer to captionGround Truth#10#30#50
Figure 9: Compression error depending on the number of used eigenbasis components. The heatmaps show the photometric $\ell_{1}$-error for 10, 30, and 50 components using $128^{2}$ Gaussian maps. Full evaluation in the supp mat.
In this section, we perform ablation of our linear GEM model in several aspects. First, we measure the memory size of such a model depending on the number of selected components compared to the size of the regressed texture. Secondly, we are interested in the compression error introduced by the projection on different amounts of principal components. Table 4 shows how the size in MB changes depending on the amount of basis and the source texture size. For instance, our smallest model weighs as little as 7MB using only 10 components of the eigenbasis. This is almost 12 times less than our smallest CNN-based model and almost 70 times less than Animatable Gaussians [38]. In contrast to neural networks, we can easily trade quality over size which is very useful in the context of different commodity devices with reduced compute capabilities. Table 4 presents how compression affects the quality of reconstruction. We evaluated $\sim 1$k frames for a single actor under a novel view. As expected using only 10 first components impacts the quality the most, however, the results are still high-quality as can be seen in Figure 9. As we previously discussed, to capture wrinkles and high-frequence details one needs to employ neural networks. Gaussian Avatars [53] offers a small size of the stored Gaussians cloud, ranging from 5MB, and 14MB without the FLAME model for $128^{2}$, $256^{2}$ Gaussians, respectively. However, the quality of reconstruction lacks wrinkle details and sharpness as can be seen in Figure 10. Note that our model does not require a FLAME model that weights an additional 90MB compared to Gaussian Avatars [53]; Table 4 presents our absolute size values.
Refer to captionGround TruthOurs #30 $128^{2}$ [53]$128^{2}$ [53]$256^{2}$
Figure 10: The quality comparison to Gaussian Avatars [53] shows better performance even though our model is similar in size to its Gaussian cloud and we do not need an additional FLAME model which weighs 90MB.
5.4 Universality of GEM
We designed GEM to be a universal method capable of distilling any 3DGS-based avatar solution [28, 38, 74, 15, 53] into a lightweight representation, as long as normalized input across training frames is provided. Our results show that a compact representation of the linear basis produces state-of-the-art results in terms of quality and speed. Figure 11 demonstrates how we can apply GEM to different primitive generators (AG [38], GA [53], Our Net). Note that to achieve wrinkle-level details, the generator itself has to produce high-quality output. As can be seen, we can apply our distillation technique to existing methods, making them light and compact. Moreover, GEM is well-suited for commodity devices, as it generates a new set of Gaussian primitives with only a single dot product needed between coefficients and the basis. We believe that this potential has great implications for downstream tasks such as holoportation, audio-driven avatars, or virtual reality.
Refer to captionGround TruthGEMGEM on AGGEM on GA
Figure 11: GEM applied on different avatar methods (AG and GA) and optimized using analysis-by-synthesis. Our method is universal and can be successfully used on both; point clouds as well as textures to distill a lightweight avatar.
6 Human Head Avatar Compression
Human avatar compression is an important topic, but it is still in its early stages and not well-explored. For neural-based representations, there are methods to compress networks, such as pruning [18], quantization [23], or knowledge distillation [20], as well as small and compact MobileNets [21]. Interestingly, in the latter context, GEM can be considered as a single-layer MLP without any activation function. Unfortunately, these methods still require an expensive forward pass and may not be well-suited for all commodity devices.
The recently introduced Gaussian Avatars by Qian [53] also represents a form of avatar compression, though not in the primitives’ space but rather in the geometry space, with Gaussians attached and deformed by triangles from a linear face model. However, this form of appearance representation is insufficient for capturing details such as wrinkles, as it rigidly adheres to FLAME rigging in the geometry space. Therefore, we advocate for different compression techniques like GEM, which can leverage more powerful representations and distill them into expressive, high-quality linear models. We hope that this project will open doors to different methods for efficiently storing and representing avatars.
7 Discussion
Despite state-of-the-art results for human face avatars, our method has still certain limitations which we would like to discuss. One of the major drawbacks of the PCA-based models is their global extent. Which ultimately means that we cannot control local changes and produce more combinations of local features beyond the standard model. Thus, further work could include incorporating a localized PCA basis [44] for better avatar control, which could potentially enable a wider range of expressions outside the training set. Another limitation is the personalization, for new subjects a new representation has to be learned from multi-view data. An interesting future avenue would be to create a statistical model across different subjects.
8 Conclusion
We have proposed Gaussian Eigen Model, a linear appearance model that can represent photo-realistic facial avatars. The simplicity of this appearance model results in massively reduced compute requirements in comparison to CNN-based avatar methods. Moreover, the lightweight representation could improve the management, sharing, and applicability of avatars. The universal distillation approach can be applied to many existing methods, making them available for compression. We demonstrate how such a representation can be used in different scenarios like self-reenactment and cross-person animation. We also show that real-time applications are possible. We believe that the proposed GEM is a stepping stone toward lightweight avatar animation on consumer-level hardware.
Gaussian Eigen Models for Human Heads
–Supplemental Materials– Wojciech Zielonka1 Timo Bolkart2 Thabo Beeler2 Justus Thies1,3
Ground Truth
$128^{2}$
$256^{2}$
$512^{2}$
#10#30#50
Figure 12: Qualitative compression quality depending on the number of basis (10, 30, 50) and resolution of the Gaussian map ($128^{2}$, $256^{2}$, $512^{2}$).
9 Compression Ablation Study
In the domain of 3D Morphable Models, principle component analysis (PCA) emerges as a cornerstone approach, instrumental in crafting the foundational framework for capturing face expressions and shapes with remarkable fidelity [3, 37]. This methodology has been adopted with notable success, not only in modeling facial features but also in extrapolating the nuances of human bodies [41, 49, 46], and even in depicting intricate hand modeling [56].
Expanding upon this foundation, GEM proposes a novel technique involving an ensemble of eigenbasis of 3D Gaussian attributes for achieving a photorealistic human head appearance. This representation exhibits significant adaptability concerning both quality and size, leveraging a fundamental trait of linear basis that proves beneficial when applied to diverse devices with varying capabilities in digital human applications. Figures 12 and 13 illustrate the qualitative and quantitative results of GEM. Notably, even under substantial compression (utilizing only ten principal components), our approach consistently yields high-quality outcomes. More examples can be found in Figures (17, 15, 18, 16)
$128^{2}$
$256^{2}$
$512^{2}$
#10#30#50
Figure 13: Compression error in PSNR (db) depending on the number of basis (10, 30, 50) and resolution of the Gaussian maps ($128^{2}$, $256^{2}$, $512^{2}$).
10 Additional Dataset Evaluation
In Figure 19, we provide further evaluation of our approach using the Multiface dataset [72]. This dataset encompasses short sequences of facial expressions, ranging from "relaxed mouth open" to "show all teeth" or "jaw open huge smile." The expressions vary widely in length and complexity, presenting a considerable challenge for analysis. It’s important to note that this dataset does not provide a parametric 3DMM; instead, it offers meshes in full correspondence. However, as mentioned in the main text, our method remains adaptable in this context. By leveraging the deformation gradient [60] to transform points from canonical space into deformed space, and assuming consistent UV parametrization of input meshes, we can successfully navigate between these spaces. As depicted in Figure 19, our network demonstrates the ability to extrapolate to novel expressions, even amidst highly challenging facial poses.
11 Broader Impact
Our project focuses on reconstructing a highly detailed human face avatar from multiview videos, enabling the extrapolation of expressions not originally captured. While our technology serves primarily constructive purposes, such as enriching telepresence or mixed reality applications, we acknowledge the potential risks of misuse. Therefore, we advocate for advancements in digital media forensics [58, 57] to aid in detecting synthetic media. Additionally, we emphasize the importance of conducting research in this area with transparency and openness, including the thorough disclosure of algorithmic methodologies, data origins, and models intended for research purposes.
Figure 14: GEM can be effectively controlled in real-time by an image-space regressor which produces coefficients projected on the linear basis of a personalized GEM avatar.
12 Future Applications & Discussion
An interesting application venue for GEM would be a combination of audio-driven methods with the appearance offered by our methods. Ng et at. [45] presented photorealistic audio-driven full-body avatars. Despite impressive results, the face region still does not fully convey expressions and lacks realism. One way of improving it would be incorporating recent progress in audio-driven geometry [66, 65, 2, 6] with a dedicated appearance model offered by GEM and our image-space regressor Figure 14.
Moreover, GEM neural network uses meshes to obtain normal maps as input to the Gaussian map regressor. However, meshes are limited by resolution and expressiveness, one way of improving on that would be to use NPHM by [13] by Giebenhain et al. and the follow-up work [61, 14, 31] to further increase the expressiveness of the model by explicitly capturing regions like hair or teeth.
Ground Truth
$128^{2}$
$256^{2}$
$512^{2}$
#10#30#50
Figure 15: Qualitative compression quality depending on the number of basis (10, 30, 50) and resolution of the Gaussian map ($128^{2}$, $256^{2}$, $512^{2}$).
$128^{2}$
$256^{2}$
$512^{2}$
#10#30#50
Figure 16: Compression error in PSNR (db) depending on the number of basis (10, 30, 50) and resolution of the Gaussian maps ($128^{2}$, $256^{2}$, $512^{2}$).
Ground Truth
$128^{2}$
$256^{2}$
$512^{2}$
#10#30#50
Figure 17: Qualitative compression quality depending on the number of basis (10, 30, 50) and resolution of the Gaussian map ($128^{2}$, $256^{2}$, $512^{2}$).
$128^{2}$
$256^{2}$
$512^{2}$
#10#30#50
Figure 18: Compression error in PSNR (db) depending on the number of basis (10, 30, 50) and resolution of the Gaussian maps ($128^{2}$, $256^{2}$, $512^{2}$).
Refer to captionGround TruthOurs NetOurs GEMGround TruthOurs NetOurs GEM
Figure 19: The Multiface dataset, introduced by Wu et al. [72], comprises actors performing scripted expressions in short segments. A notable challenge arises due to the occurrence of several expressions, like "show all teeth," appearing only once in the dataset. This poses a difficulty during testing, particularly when the network is required to extrapolate. Here we showcase the outcomes of the test sequences to illustrate the effectiveness of our CNN-based network in capturing diverse and challenging facial poses, demonstrating its robustness despite the inherent complexity of the dataset.
Refer to captionGround TruthOurs GEMAG [38]GA [53]PointAvatar [82]AvatarMAV[75]
Figure 20: Additional baselines PointAvatar (PSNR: 25.8, SSIM: 0.893 LPIPS: 0.097) and AvatarMAV (PSNR: 29.5, SSIM: 0.913, LPIPS: 0.152) evaluated on the novel-view sequences.